OpenStudy (anonymous):

write an equation in slope-intercept form of the line parallel to y=2x-6;containing (8,1)

OpenStudy (anonymous):

Ok. Slope-intercept form is the following: y=mx +b, where m is the slope and b is the y-intercept. You want a parallel line to y=2x-6, so your slope has to be 2 (since two lines are parallel if and only if they have the same slopes). So, by now we have y=2x+b : We have to find what b is. Which is easy. They give you a point in the line, (8, 1). You substite these values for x and y respectively, then solve for b. (1)=2(8)+b 1= 16+b -15=b Then, your line is the following: y=2x-15

OpenStudy (anonymous):

nice! thx!